OpenStudy (anonymous):

Tangent Line Approximation Problem: If a right triangle has legs 6 and 8, its hypotenuse is 10. The triangle will be inscribed within a circle with area 25pi . (The hypotenuse will be the diameter of the circle.) A. Suppose one leg of the triangle is known to be exactly 6, but the other leg is known to be 8 with an error of +/-h . What are x, f(x), and a in this problem?

OpenStudy (anonymous):

B. Use a tangent line approximation to estimate the area of the circumscribed circle. C. Now consider the sphere that just contains the triangle (so the hypotenuse is the diameter of the sphere). Use a tangent line approximation to estimate the volume of this sphere This is free response; can someone help me get started? .

OpenStudy (amistre64):

we get options?

OpenStudy (anonymous):

I'm sorry, what do you mean?

OpenStudy (amistre64):



OpenStudy (amistre64):

at means it looks like your taking a test and wanting us to pick the right option :)

OpenStudy (anonymous):

Oh! No, those are separate questions. This is a homework practice assignment for Tangent Line Approximation in Calc 1.

OpenStudy (amistre64):



OpenStudy (amistre64):

then im prolly butchering this thing up :)

OpenStudy (anonymous):

That looks like my picture! But what does the text say?

OpenStudy (amistre64):

not the center, im making note of that even tho it looks centery like

OpenStudy (amistre64):

centernal?

OpenStudy (anonymous):

I have no idea :P The diameter is 10 though, correct?

OpenStudy (amistre64):

hyp is diameter, oy; its like reading a novel by steinbeck

OpenStudy (amistre64):

what is an "a"? is there a pic that goes with this?

OpenStudy (anonymous):

Oh! The equation for tangent line approximation is f(x) is approximately equal to f(a)+f'(a)(x-a)

OpenStudy (anonymous):

If we were just trying to change the area of the triangle, I believe, because 8 is +/- h, 8 would be a and 8+h or 8-h would be x.

OpenStudy (amistre64):

the diameter is the hyp; and the hyp depend on 8+h or 8-h 6^2 + (8-h)^2 = lesser margin ^2 6^2 + (8+h)^2 = higher margin ^2

OpenStudy (anonymous):

Wait! The area of the circle is \[\pi r^2\] The r is dependent on \[6^2 + (8\pm h)^2\]

OpenStudy (amistre64):

correct

OpenStudy (anonymous):

So my function that I'll be using will be a synthesis of these two. Any suggestions of a function to represent this?

OpenStudy (anonymous):

The area of the circle, I mean.

OpenStudy (amistre64):

replace r^2 by its substitution

OpenStudy (amistre64):

pi(36+(8+-h)^2)

OpenStudy (anonymous):

Does the portion in parentheses need to be squared? It's pi r^2, right?

OpenStudy (anonymous):

Oh. Nevermind.

OpenStudy (amistre64):

you see it right :)

OpenStudy (anonymous):

Yes! I do! Thanks so much for your help. Okay, moving onto the sphere portion...The same thing applies with the radius being substituted except it will be substituted into the volume of a sphere equation?

OpenStudy (amistre64):

yes, since the volume of a sphere is dependant on "r" 4/3 pi r^3 is the volume so make a neccesary adjustment for the extra "r" r = sqrt(36 + (8+-h)^2 )

OpenStudy (anonymous):

Thank you so much! I should be able to continue.

OpenStudy (amistre64):

good luck :)